OpenStudy (anonymous):

solve the system using elemination 4x+2y=8 2x-2y=4

OpenStudy (solomonzelman):

4x+2y=8 2x-2y=4 ----> times (-2) ----> -4x+4y=-8 4x +2y = 8 -4x +4y = -8 ----------------- ADD THEM.

OpenStudy (anonymous):

how!

OpenStudy (solomonzelman):

4x +2y = 8 -4x +4y = -8 ----------------- (4-4)x + (2+4)y = (8-8) 0x + 6y = 0 we got rid of x, 6y=0 y=0 can you find x?